OpenStudy (alexh107):

Use the square root method to find the solutions to 3x^2+2=0

OpenStudy (alexh107):

Can someone tell me where I'm going wrong: \[3x^2+2=0\] \[3x^2=-2\] \[x^2=-2/3\] \[\sqrt{x^2}=\sqrt{-2/3}\] \[x=+-i \sqrt{2/3}\]

Awolflover1 (awolflover1):

everything is correct but the last one is positive...

OpenStudy (alexh107):

The answer in the book says it should be \[+-i \sqrt{6}/3\]

Awolflover1 (awolflover1):

hmm...